lim [x^2 sin(1/x)/(sinx) =1

x to 0

True or false with full explanation


1
Expert's answer
2021-07-06T10:04:04-0400

"L=\displaystyle \lim_{x\to 0} \frac{x^2 sin(\frac{1}{x})}{sin \ x}=\displaystyle \lim_{x\to 0} \frac{x sin(\frac{1}{x})}{\frac{sin \ x}{x}}" "=" "\displaystyle \lim_{x\to 0} \frac{\frac{sin(1/x)}{1/x}}{\frac{sin \ x}{x}}"

"=" "\frac{\displaystyle \lim_{x\to 0}{\frac{sin(1/x)}{1/x}}}{\displaystyle \lim_{x\to 0}\frac{sin \ x}{x}}"

Now, "\displaystyle \lim_{x\to 0}\frac{sin \ x}{x}=1"

So , "L=\displaystyle \lim_{x\to 0}{\frac{sin(1/x)}{1/x}}=\displaystyle \lim_{x\to 0} x\ sin (1/x)"

and "\displaystyle \lim_{x\to 0} x\ sin\ \theta =0 \ \forall \ \theta \isin \ R"

"L=\displaystyle \lim_{x\to 0} x\ sin (1/x)=0"

Hence, the statement "L=\displaystyle \lim_{x\to 0} \frac{x^2 sin(\frac{1}{x})}{sin \ x}=1" is wrong.